what is 6+y=8 and is it linear?
Answer by ikleyn(52915)

.

6 + y = 8 is an equation.

It is a linear equation, since it includes variable y in degree 1
and does not include variable y in other degrees.

Its solution is y = 8 - 6 = 2.

------------------

If you want to learn on how to solve simple and simplest linear equations,  read the lesson
    - HOW TO solve a linear equation
in this site.

Consider this lesson as your textbook,  handbook,  tutorials and  (free of charge)  home teacher.

Happy learning  (!)